# Sliding Speed Influences Electrovibration-Induced Finger Friction Dynamics on Touchscreens

### v1

@@ Abstract
Electrovibration technology can render tactile textures on capacitive touchscreens by modulating friction between the finger and the screen through electrostatic attraction force generated by applying an alternating voltage signal to the screen. This signal should be carefully calibrated for realistic and robust texture rendering. However, this process is challenging due to variations in sliding speed, applied force, and individual skin mechanics, which affect friction in complex and unpredictable ways. Here, we investigate how exploration conditions affect electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this process. Ten participants slid their index fingers across an electrovibration-enabled touchscreen at five sliding speeds ($20\sim 100$ mm/s) and applied force levels ($0.2\sim 0.6$ N) while we measured contact forces and skin accelerations. The touchscreen was excited with amplitude-modulated voltage signals across frequencies relevant to touch. We modeled the finger-touchscreen friction response as a first-order system and the skin mechanics as a mass-spring-damper system. Our results showed that the sliding speed influenced the cutoff frequency of the friction response as well as the moving mass and stiffness of the finger for the tested exploration ranges. Specifically, for every 1 mm/s increase in speed, the cutoff frequency, the finger moving mass, and stiffness increased by 13.8 Hz, $3.23\times 10^{-5}$ kg, and 4.04 N/m, respectively. Further correlation analysis revealed that finger stiffness affected the cutoff frequency more than the moving mass. Finally, we developed a practical model for electrovibration-induced finger friction on touchscreens that accounts for sliding speed variations, paving the way for delivering consistent haptic feedback through electrovibration.
@@
Surface haptics , Electrovibration , Skin mechanics , System identification , Texture rendering
journal: arXiv
@@
A model was proposed to compensate for speed-based variations in electrovibration friction
@@ 1 Introduction
Electrovibration is one such technology that can modulate friction between a touchscreen and a user’s finger via electrostatic forces [2]. When a finger interacts with the touchscreen, it functions as a parallel plate capacitor, where the screen’s non-conductive coating, the outermost layer of the skin (stratum corneum), and an air gap between the skin and the touchscreen (caused by the finger ridges and touchscreen’s rough surface) act as the dielectric layers [3, 4, 5]. The touchscreen’s conductive layer beneath the non-conductive coating and the skin’s inner tissues serve as two parallel conductors. Applying an alternating high-voltage signal to the conductive layer of the touchscreen generates an electrostatic force between the finger and the touchscreen, pulling the skin toward the surface in response to the input voltage signal. When the finger moves, this attraction and release of the skin alter the friction between the finger and the screen, creating the sensation of texture on an otherwise smooth glass surface. By using recorded contact forces from real-world finger-surface interactions as input signals, electrovibration displays can replicate virtual texture sensations that closely match their physical counterparts [6, 7]. Electrovibration outperforms other surface haptic technologies for its texture rendering capability, as it has a wide bandwidth, generates localized skin vibrations across the touchscreen, and the produced sensations are independent of the screen size [8].
@@
In this paper, we investigate how exploration conditions influence the friction between a finger and an electrovibration display, and the role that skin mechanics play in this interaction. We measured the generated friction forces and skin vibrations when 10 participants explored an electrovibration display under controlled conditions, varying both applied force (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N) and sliding speed (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s). The display was driven with 15 logarithmically spaced input frequencies (30 Hz $\sim$ 2 kHz) modulated using the technique described by Shultz et al. [18]. We then identified and modeled the electromechanical behavior between input voltage and measured friction force to analyze how the model parameters change as a function of applied forces and speeds. Additionally, since skin biomechanical properties vary among individuals, we analyzed the model’s variability across 10 participants. Finally, we investigated the relationship between the skin’s mechanical properties and the corresponding electro-mechanical response. We believe that this study offers valuable new perspectives for developing more realistic texture interactions on touchscreens.
@@ 2.1 Electrovibration
Electrovibration modulates friction between a capacitive touchscreen and a moving finger using electrostatic forces. The resulting friction force, $F_{f}$, is influenced by both the applied normal force exerted by the finger, $F_{n}$, and the electrostatic force, $F_{e}$. The relation $F_{f}=\mu(F_{n}+F_{e})$ describes the friction force, where $\mu$ represents the friction coefficient. The $F_{e}$ depends on the applied input voltage, $V_{i}$, and is given by the equation [3]:
@@
As mentioned in Section 1, eliminating this nonlinearity and any charge leakage is necessary for minimizing the discrepancy between the input voltage and output friction for realistic texture rendering. Due to these physical phenomena, the input voltage spectrum is attenuated at low frequencies and spreads, causing the output friction spectrum to mismatch with the input voltage spectrum. As humans are sensitive to tactile frequencies up to 1 kHz [30], any output friction distortions below 1 kHz may cause a different texture feel than intended, reducing the virtual texture realism.
@@ 2.2 Experimental setup
The touchscreen was placed on a force sensor (Nano43 Titanium, ATI) using double-sided tape (Tesa PRO) to measure the generated forces in three dimensions. We mounted the force sensor on a large aluminum block with screws. This entire assembly was placed on an aluminum optical breadboard table (MB60120/M, Thorlabs), supported by a sturdy frame (PFM52502, Thorlabs). A monitor (UltraSharp 24, Dell) displayed graphical information to the participants. The data acquisition card collected data from the force sensor at a sampling rate of 20 kHz. A 2D infrared position sensor (NNAMC1580PCEV, Neonode) tracked the finger’s position with a resolution of 0.1 mm and a sampling rate of 60 Hz. The sensor was mounted on a 3D-printed support, ensuring its active plane was positioned above the touchscreen. We enclosed the touchscreen in a box with a 70$\times$80 mm cutout for finger exploration. We also measured the three-dimensional skin vibrations via an accelerometer (ADXL358, Analog Devices) attached to the participants’ fingernails via an accelerometer wax (Model 080A109, PCB Piezotronics). The accelerometer data was also collected through the same data acquisition card with a sampling rate of 20 kHz.
@@ 2.3 Procedure
The aim of the experiments was to record the finger interaction forces and accelerations while the participants explored the electrostatically actuated touchscreen at a 60 contact angle. The cut-out of the enclosure box around the touchscreen ensured that the finger always maintained a constant contact angle with the touchscreen (see Figure 2). During the experiments, a participant moved the index finger of their dominant hand on the touchscreen horizontally (left to right) at five different sliding speeds (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s) while applying five different normal forces (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N). A custom graphical user interface (GUI) guided the participants by displaying two visual square cues: the top square indicated the reference speed, while the bottom square tracked the user’s finger movement (Figure 2). The tracking square changed color to indicate the applied force: green for the target range, yellow for below the target, and red for above the target. If the participant’s mean, minimum, and maximum exploration force and speed deviated by ±$\pm 10\%$ of the target range for more than one second, then the trial was discarded and repeated.
@@
Each trial lasted 17 seconds and consisted of two phases. During the first 7 seconds, the participant adjusted their finger exploration to the experimental conditions; however, their interaction data was not recorded during this period. In the remaining 10 seconds, we recorded lateral force and skin vibration data. During this period, the participants swiped the touchscreen from left to right six times. After each trial, the process was repeated for subsequent trials.
@@ 2.4 Data processing
During the experiments, we recorded the contact forces, finger accelerations, and the input voltage signal after modulation for analysis, as shown in Figure 3(a). The finger’s active exploration was, in general, not exactly a straight line; this generated lateral forces, which were primarily recorded by the force sensor’s y-axis and partially along the x-axis. To capture the generated friction forces accurately, we transformed the two-dimensional lateral force data into a single dimension in the frequency domain [33], extracting the friction signal as shown in Figure 3(b). Similarly, the accelerometer’s x-axis primarily recorded the lateral finger vibrations along the finger exploration direction. To align the accelerometer axes with the force sensor axes, we multiplied two rotation matrices: one along the x-axis at $155^{\circ}$ and another along the y-axis at $180^{\circ}$, accounting for the inverse orientation of the accelerometer on the y-axis and the finger exploration angle of $60^{\circ}$. After multiplying the raw sensor readings by $9.8m/s^{2}$ for conversion to a true acceleration signal, we integrated the signals in time to obtain the velocity data, as shown in Figure 3(c).
@@
After extracting the samples for the input amplitude-modulated voltage signals, $V_{i}$, for different finger sweeps, we demodulated them to get the original message signals back to model the input message signal, $V_{m}$, to output friction, $F_{f}$, response. To do this, we applied the Fourier transform on the input amplitude-modulated signal, $V_{i}$, and then we calculated the magnitude and phase of the upper envelope, $\omega_{c}+\omega_{m}$, of the amplitude-modulated signal. Using the magnitude and phase value, we reconstructed the sinusoidal message signal. We repeated the same process for all the sweep data, frequencies, exploration conditions, and participants.
@@ 2.5 Dynamic modeling of electrovibration-induced finger friction on a touchscreen
Our goal was to investigate how exploration speed and force influence the electrovibration-induced finger friction on a touchscreen. To achieve this, we modeled the relationship between the input message signal, $V_{m}$, and output friction, $F_{f}$, while assessing the impact of exploration conditions on the model parameters. Given the nonlinear behavior of the fingertip when interacting with a surface [34], we sought to simplify the electrovibration-induced finger friction model by conditionally representing it as a linear system. To check this, we identified the transfer function, $H_{f-e}(s)$, by dividing the Fourier transform of the output friction, $F_{f}$, by the Fourier transform of the input message signal, $V_{m}$, for a particular exploration force and speed. We repeated this process for every finger sweep made by the participant and every test frequency to construct a Bode plot as shown in Figure 4; the input-output dynamics behave like a first-order low-pass up to 750 Hz and then show two resonance peaks at 860 Hz and 1740 Hz. We verified that the two resonance peaks are due to the experimental setup’s hardware; please refer to B for more details. We used the fmincon algorithm in MATLAB to optimize the transfer equation based on the discrete transfer function points found earlier. Then, we removed the setup’s dynamics as given by Equation (23) from the fitted transfer function model. As a result, we found a first-order response for the input message signal, $V_{m}$, to output friction, $F_{f}$, response, as shown in Equation (3).
@@
This first-order behavior aligns with previous findings by Meyer et al. [17], who demonstrated that the message signal of an amplitude-modulated voltage signal and the output friction, $F_{f}$, in finger-electrovibration interaction followed a first-order low-pass response at an arbitrary exploratory force and speed. They modeled the behavior using the device’s quality factor, $Q$, and carrier frequency, $\omega_{c}$, as $H_{f-e}=\frac{Q}{1+{2Qs}/{\omega_{c}}}$. However, to better represent the input message signal, $V_{m}$, to output friction, $F_{f}$, response variation caused by exploration speed and force, we opted for a generic model given by Equation (3). Figure 4 shows an input message signal, $V_{m}$, to output friction, $F_{f}$, response computed for one of our participants at 40 mm/s exploration speed and 0.2 N normal force. The corresponding estimated model parameters are $K=0.0147$ (N/V), and $\omega_{o}=961$ Hz.
@@ 2.6 Dynamic modeling of finger skin mechanical behavior
Additionally, we modeled finger vibrations caused by the friction signal to investigate how the finger’s mechanical behavior influences electrovibration-induced finger friction on a touchscreen. Here, we considered the measured electrovibration-induced finger friction force, $F_{f}$, as input and finger vibration velocity, $v$, as output. As the finger exhibits a nonlinear behavior [35, 34, 36], we wanted to know if the input friction force, $F_{f}$, to output finger vibration velocity, $v$, can be represented by a linear system. To do this, we divided the Fourier transform of the finger vibration signal by the Fourier transform of the friction signal to obtain the transfer function, $H_{f}$. We repeated the process for every finger sweep and frequency for a particular exploration condition, similar to finger-electrovibration modeling. After plotting the transfer function points using a Bode plot, see Figure 5, we observed that the system showed a second-order response. Using the fmincon algorithm in Matlab, we modeled the system based on the commonly used mass, spring, and damper system in the literature [15, 16, 7]:
@@
Here, $v(s)$ and $F_{f}(s)$ are the Fourier transform of the finger vibration as velocity signal and friction generated due to finger electrovibration interaction, $H_{f}(s)$ is the transfer function between them, $m$ is finger-moving mass, $b$ is finger damping, and $k$ is finger stiffness. The moving mass of the finger represents the effective mass of the subcutaneous finger pad, while the damping and stiffness correspond to the viscosity and elasticity of the fingertip pulp [15]. Figure 5 shows an example finger mechanical response obtained when the display was explored with 40 mm/s sliding speed and 0.2 N applied force. The finger mechanical model parameters for the response are $m=0.0015$ Kg, $b=1.3$ Ns/m, $k=444$ N/m. Our findings matched the mechanical response, their derived mass, spring, and damping values reported in the previous studies [15, 16, 7], even though there are other techniques to excite and measure the finger response.

### v2

@@ Abstract
Electrovibration technology enables tactile texture rendering on capacitive touchscreens by modulating friction between the finger and the screen through electrostatic attraction forces, generated by applying an alternating voltage signal to the screen. Accurate signal calibration is essential for robust texture rendering but remains challenging due to variations in sliding speed, applied force, and individual skin mechanics, all of which unpredictably affect frictional behavior. Here, we investigate how exploration conditions affect electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this process. Ten participants slid their index fingers across an electrovibration-enabled touchscreen at five sliding speeds ($20\sim 100$ mm/s) and applied force levels ($0.2\sim 0.6$ N). Contact forces and skin accelerations were measured while amplitude modulated voltage signals spanning the tactile frequency range were applied to the screen. We modeled the finger-touchscreen friction response as a first-order system and the skin mechanics as a mass-spring-damper system. Results showed that sliding speed influenced the friction response’s cutoff frequency, along with the estimated finger moving mass and stiffness. For every 1 mm/s increase in speed, the cutoff frequency, the finger moving mass, and stiffness increased by 13.8 Hz, $3.23\times 10^{-5}$ kg, and 4.04 N/m, respectively. Correlation analysis revealed that finger stiffness had a greater impact on the cutoff frequency than moving mass. Notably, we observed a substantial inter-participant variability in both finger-display interaction and skin mechanics parameters. Finally, we developed a speed-dependent friction model to support consistent and perceptually stable electrovibration-based haptic feedback across varying user conditions.
@@
Biotribology , Skin mechanics , Finger-surface friction , Surface haptics , Electrovibration , Texture rendering
journal: Tribology International
@@
Electrovibration-induced finger friction dynamics vary across participants
A model is proposed to compensate for speed-based variations in electrovibration friction
@@ 1 Introduction
Electrovibration is one such technology that can modulate friction between a touchscreen and a user’s finger via electrostatic forces [2]. When a finger interacts with the touchscreen, it functions as a parallel plate capacitor, where the screen’s non-conductive coating, the outermost layer of the skin (stratum corneum), and an air gap between the skin and the touchscreen (caused by the finger ridges and touchscreen’s rough surface) act as the dielectric layers [3, 4, 5]. The touchscreen’s conductive layer beneath the non-conductive coating and the skin’s inner tissues serve as two parallel conductors. Applying an alternating high-voltage signal to the conductive layer of the touchscreen generates an electrostatic force between the finger and the touchscreen, pulling the skin toward the surface in response to the input voltage signal. Although this force is very weak when the finger is static [6], it modulates friction during finger movement by repeatedly attracting and releasing the skin, thereby creating the sensation of texture on an otherwise smooth glass surface. By using recorded contact forces from real-world finger-surface interactions as input signals, electrovibration displays can replicate virtual texture sensations that closely match their physical counterparts [7, 8]. Electrovibration outperforms other surface haptic technologies for its texture rendering capability, as it has a wide bandwidth, generates localized skin vibrations across the touchscreen, and the produced sensations are independent of the screen size [9].
@@
In this paper, we investigate how exploration conditions influence the friction between a finger and an electrovibration display, as well as the role of skin mechanics in this interaction, with particular emphasis on their implications for texture rendering. We measured the generated friction forces and skin vibrations when 10 participants explored an electrovibration display under controlled conditions, varying both applied force (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N) and sliding speed (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s). The display was driven with 15 logarithmically spaced input frequencies (30 Hz $\sim$ 2 kHz) modulated using the technique described by Shultz et al. [18]. We then identified and modeled the relationship between input voltage and measured friction force, as well as the mechanical response of the finger to the generated friction. For both models, we analyzed how the model parameters varied as a function of applied force and sliding speed. Additionally, since skin biomechanical properties vary among individuals, we examined model variability across participants. Finally, we investigated the relationship between the measured skin mechanical properties and the corresponding model parameters of the electrovibration-induced finger friction response.
@@ 2.1 Electrovibration
Electrovibration modulates friction between a capacitive touchscreen and a moving finger using electrostatic forces. The resulting friction force, $F_{f}$, is influenced by both the applied normal force exerted by the finger, $F_{n}$, and the electrostatic force, $F_{e}$. The relation, $F_{f}=\mu(F_{n}+F_{e})$, describes the friction force, where $\mu$ represents the friction coefficient. The $F_{e}$ depends on the applied input voltage, $V_{i}$, and is given by the equation [3]:
@@
As mentioned earlier, eliminating this nonlinearity and any charge leakage is necessary for minimizing the discrepancy between the input voltage and output friction for realistic texture rendering. Due to these physical phenomena, the input voltage spectrum is attenuated at low frequencies and spreads, causing the output friction spectrum to mismatch with the input voltage spectrum. As humans are sensitive to tactile frequencies up to 1 kHz [30], any output friction distortions below 1 kHz may cause a different texture feel than intended, reducing the virtual texture realism.
@@ 2.2 Experimental setup
[…]
@@ 2.3 Procedure
The aim of the experiments was to record the finger interaction forces and accelerations while the participants explored the electrostatically actuated touchscreen at a 60 contact angle. This specific angle was chosen based on prior research indicating that a 60 contact angle naturally occurs during texture discrimination tasks [33, 34]. Furthermore, higher contact angles have been shown to increase finger stiffness, thereby enhancing the transmission of vibratory cues to the fingertip [35, 36]. This improvement in vibration transmission contributes to better texture perception, which is relevant for texture rendering. The cut-out of the enclosure box around the touchscreen ensured that the finger always maintained a constant contact angle with the touchscreen (see Figure 2).
During the experiments, a participant moved the index finger of their dominant hand on the touchscreen horizontally (left to right) at five different sliding speeds (20 mm/s, 40 mm/s, 60 mm/s, 80 mm/s, and 100 mm/s) while applying five different normal forces (0.2 N, 0.3 N, 0.4 N, 0.5 N, and 0.6 N). These force and speed values were selected based on previous studies that reported typical ranges observed during natural active touch exploration – approximately 0.15 N to 0.64 N for normal force [33, 36, 37], and 10 mm/s to 100 mm/s for sliding speed [34, 23, 37]. Horizontal movement was specifically chosen because the finger predominantly exhibits rolling motion in this direction, which is associated with a lower coefficient of friction and reduces the likelihood of stick-slip behavior [38].
A custom graphical user interface (GUI) guided the participants by displaying two visual square cues: the top square indicated the reference speed, while the bottom square tracked the user’s finger movement (Figure 2). The tracking square changed color to indicate the applied force: green for the target range, yellow for below the target, and red for above the target. If the participant’s mean, minimum, and maximum exploration force and speed deviated by $\pm 10\%$ and $\pm 25\%$ of the target range for more than one second, then the trial was discarded and repeated. More details regarding this procedure can be found in C.
@@
Each trial lasted 17 seconds and consisted of two phases. During the initial 7 seconds, the participant adjusted their finger movements to match the experimental conditions; however, no interaction data was recorded during this period. In the subsequent 10 seconds, we measured the friction between the fingertip and the touchscreen using a force sensor mounted beneath the surface, as participants slid their finger at the instructed speed and applied normal force. This measurement captured both the baseline friction with the smooth glass and the additional friction induced by electrovibration. Concurrently, we recorded fingertip skin vibrations using an accelerometer attached to the participant’s fingernail. Participants typically completed more than six swipes in 10 seconds—–approximately three lateral and three medial. At the lowest speed (20 mm/s), each 70 mm swipe took around 2.35 s, reducing the average to three lateral and two medial swipes. The procedure was repeated for each subsequent trial.
@@ 2.4 Data processing
During the experiments, we recorded the contact forces, finger accelerations, and the input voltage signal after modulation for analysis, as shown in Figure 3(a). The finger’s active exploration generally did not follow a perfectly straight path (see Figure 14). The resulting lateral forces were primarily captured along the force sensor’s y-axis, with partial contributions along the x-axis. To capture the generated friction forces accurately, we transformed the two-dimensional lateral force data into a single dimension in the frequency domain [39], extracting the friction signal as shown in Figure 3(b). Similarly, the accelerometer captured finger vibrations caused by friction along its three axes, with the x-axis primarily recording lateral vibrations in the direction of finger exploration and residuals along the y and z axes. To align the accelerometer axes with the force sensor axes, we multiplied two rotation matrices: one along the x-axis at $155^{\circ}$ and another along the y-axis at $180^{\circ}$, accounting for the inverse orientation of the accelerometer on the y-axis and the finger exploration angle of $60^{\circ}$. After converting raw sensor readings to true acceleration values by multiplying them by $9.8m/s^{2}$, we applied the DFT321 algorithm [39] to reduce the 3D signals and obtain the resultant acceleration along the exploration direction. This signal was then integrated over time to obtain the velocity data, as shown in Figure 3(c).
@@
After extracting the samples for the input amplitude-modulated voltage signals, $V_{i}$, for different finger sweeps, we demodulated them to get the original message signals back to model the input message signal, $V_{m}$, to output the friction, $F_{f}$, response. To do this, we applied the Fourier transform on the input amplitude-modulated signal, $V_{i}$, and then we calculated the magnitude and phase of the upper envelope, $\omega_{c}+\omega_{m}$, of the amplitude-modulated signal. Using the magnitude and phase value, we reconstructed the sinusoidal message signal. We repeated the same process for all the sweep data, frequencies, exploration conditions, and participants.
@@ 2.5 Dynamic modeling of electrovibration-induced finger friction on a touchscreen
Our goal was to investigate how exploration speed and force influence the electrovibration-induced finger friction on a touchscreen, with the primary aim of understanding this behavior in the context of texture rendering. To achieve this, we modeled the relationship between the input message signal, $V_{m}$, and output friction, $F_{f}$, while assessing the impact of exploration conditions on the model parameters. Given the nonlinear behavior of the fingertip when interacting with a surface [40], we sought to simplify the electrovibration-induced finger friction model by conditionally representing it as a linear system. To check this, we identified the transfer function, $H_{f-e}(s)$, by dividing the Fourier transform of the output friction, $F_{f}$, by the Fourier transform of the input message signal, $V_{m}$, for a particular exploration force and speed. We repeated this process for every finger sweep made by the participant and every test frequency to construct a Bode plot as shown in Figure 4; the input-output dynamics showed a 20 dB/ decade roll off rate matching a first-order low-pass up to 750 Hz and then showed two resonance peaks at 860 Hz and 1740 Hz. We verified that the two resonance peaks are due to the experimental setup’s hardware; please refer to B for more details. We used the fmincon algorithm in MATLAB to optimize the transfer equation based on the discrete transfer function points found earlier. Then, we removed the setup’s dynamics as given by Equation (24) from the fitted transfer function model. As a result, we found a first-order response for the input message signal, $V_{m}$, to output friction, $F_{f}$, response, as shown in Equation (3).
@@
This first-order behavior aligns with previous findings by Meyer et al. [17], who demonstrated that the message signal of an amplitude-modulated voltage signal and the resulting friction force, $F_{f}$, exhibited a first-order low-pass response at an arbitrary exploratory force and speed. They modeled the behavior using the device’s quality factor, $Q$, and carrier frequency, $\omega_{c}$, as $H_{f-e}=\frac{Q}{1+{2Qs}/{\omega_{c}}}$. However, to better represent the input message signal, $V_{m}$, to output friction, $F_{f}$, response variation caused by exploration speed and force, we opted for a generic model given by Equation (3). Figure 4 shows an input message signal, $V_{m}$, to output friction, $F_{f}$, response computed for one of our participants at 40 mm/s exploration speed and 0.2 N normal force. The corresponding estimated model parameters are $K=0.0147$ (N/V), and $\omega_{o}=961$ Hz.
@@
While higher-order models can also be used to capture the more complex dynamics of electrovibration-induced finger friction, we opted for a low-order model. Higher-order models risk overcompensation during texture rendering, potentially leading to instability and unnecessary complexity. In contrast, low-order models offer a more robust and practical solution for real-time rendering applications.
@@ 2.6 Dynamic modeling of finger skin mechanical behavior
Additionally, we modeled finger vibrations caused by the friction signal to investigate how the finger’s mechanical behavior influences electrovibration-induced finger friction on a touchscreen. Here, we considered the measured electrovibration-induced finger friction force, $F_{f}$, as input and finger vibration velocity, $v$, as output. As the finger exhibits a nonlinear behavior [41, 40, 42], we wanted to know if the input friction force, $F_{f}$, to output finger vibration velocity, $v$, can be represented by a linear system. With this aim, we divided the Fourier transform of the finger vibration signal by the Fourier transform of the friction signal to obtain the transfer function, $H_{f}$. We repeated the process for every finger sweep and frequency for a particular exploration condition, similar to electrovibration-induced finger friction modeling. After plotting the transfer function points using a Bode plot, see Figure 5, we observed that the system showed a second-order response. Then, using the fmincon algorithm in Matlab, we modeled the system based on the commonly used mass, spring, and damper system in the literature [15, 16, 8]:
@@
Here, $v(s)$ and $F_{f}(s)$ are the Fourier transform of the finger vibration as velocity signal and friction generated due to finger electrovibration interaction, $H_{f}(s)$ is the transfer function between them, $m$ is finger-moving mass, $b$ is finger damping, and $k$ is finger stiffness. The moving mass of the finger represents the effective mass of the subcutaneous finger pad, while the damping and stiffness correspond to the viscosity and elasticity of the fingertip pulp [15]. Figure 5 shows an example finger mechanical response obtained when the display was explored with 40 mm/s sliding speed and 0.2 N applied force. The finger mechanical model parameters for the response are $m=0.0015$ Kg, $b=1.3$ Ns/m, $k=444$ N/m. Our findings matched the mechanical response, their derived mass, spring, and damping values reported in the previous studies [15, 16, 8], despite differences in their excitation and measurement techniques.